OpenStudy (anonymous):

Am I doing this right? find dy/dx by implicit differentiation. 2sinxcosy

OpenStudy (anonymous):

sinxcosy=1/2 cosxcosy+sinx(-siny\[\left(\begin{matrix}dy \\ dx\end{matrix}\right)\])=1/2 -siny(dy/dx) = (1/2-cosxcosy)/sinx dy/dx = (1-cosxcosy)/(-2sinysinx) right?